# Accurate, Robust and Privacy-Preserving Brain-Computer Interface Decoding

### v1

@@ Abstract
An electroencephalogram (EEG) based brain-computer interface (BCI) enables direct communication between the brain and external devices. However, EEG-based BCIs face at least three major challenges in real-world applications: data scarcity and individual differences, adversarial vulnerability, and data privacy. While previous studies have addressed one or two of these issues, simultaneous accommodation of all three challenges remains challenging and unexplored. This paper fills this gap, by proposing an Augmented Robustness Ensemble (ARE) algorithm and integrating it into three privacy protection scenarios (centralized source-free transfer, federated source-free transfer, and source data perturbation), achieving simultaneously accurate decoding, adversarial robustness, and privacy protection of EEG-based BCIs. Experiments on three public EEG datasets demonstrated that our proposed approach outperformed over 10 classic and state-of-the-art approaches in both accuracy and robustness in all three privacy-preserving scenarios, even outperforming state-of-the-art transfer learning approaches that do not consider privacy protection at all. This is the first time that three major challenges in EEG-based BCIs can be addressed simultaneously, significantly improving the practicalness of EEG decoding in real-world BCIs.
@@ I Introduction
A typical closed-loop EEG-based BCI system includes four essential parts: signal acquisition, signal processing, machine learning, and a controller, as shown in Fig. 1.
@@
However, to our knowledge, no study so far can address all three challenges simultaneously. This paper fills this gap. We demonstrate for the first time that it is feasible to achieve simultaneously accuracy, robust and privacy-preserving decoding of EEG-based BCIs through a single classifier. More specifically, we first propose Augmented Robustness Ensemble (ARE), which leverages data alignment, data augmentation, adversarial training, and ensemble learning to improve the classification accuracy and robustness, and then integrate it into three privacy-preserving transfer learning scenarios:
@@
    We propose ARE, which leverages data alignment, data augmentation, adversarial training, and ensemble learning to improve simultaneously the classification accuracy and adversarial robustness.
@@
    We integrate ARE with privacy-preserving transfer learning, achieving for the first time simultaneous accurate decoding, adversarial robustness, and privacy protection.
@@ II-C Privacy Protection
III Accurate, Robust, and Privacy-Preserving BCI Decoding
This section introduces our proposed approach for accurate, robust, and privacy-preserving EEG-Based BCIs, including the ARE algorithm and three privacy-preserving scenarios.
@@ III-B Overall Framework
In privacy-protection scenarios, the task model $\emph{C}_{\bm{\theta}}$ trained in the source domain or perturbed privacy-preserving source data $\tilde{\mathcal{D}}_{S}$ can be used to facilitate the training in the target domain using our proposed ARE, as shown in Fig. 2 and Algorithm 1.
@@
Algorithm 1 Algorithm for accurate, robust and privacy-preserving BCI decoding.
@@ III-C1 Centralized Source-Free Transfer Learning
In this scenario, the source users trust each other, so they can share raw EEG data among themselves, but not with the target user. EEG data from multiple source users are aggregated to train a single source model, which is then utilized in transfer learning to train the target model, as illustrated in Fig. 3. As the target user cannot access the source domain data, the source data privacy is protected.
@@ III-C2 Federated Source-Free Transfer Learning
In this more strict privacy-protection scenario, no source user wants to share his/her raw EEG data with other source users and the target user. A central server, without access to the source users’ private EEG data, maintains a global model and distributes it to each source user for updates. Each source user refines the global model parameters using his/her own EEG data and sends these updates back to the server for aggregation, as illustrated in Fig. 4. Our recently proposed federated learning algorithm [44], which achieves superior performance compared to centralized training, is utilized in this paper.
@@ III-C3 Source Data Perturbation
A user-specific perturbation can be applied to each source domain to safeguard source user privacy while still leveraging the source data for target domain classifier training, as illustrated in Fig. 5. This perturbation makes the user identity information unlearnable while preserving the BCI task related information. Our recently proposed privacy-preserving perturbation approach [26] is utilized in this paper.
@@
III-D Augmented Robustness Ensemble
We propose ARE to train an accurate and robust EEG classifier for the target user, based on the target domain calibration set $\mathcal{D}_{T}=\left\{\left(X_{T,i},y_{T,i}\right)\right\}_{i=1}^{N_{T}}$. ARE has the following four steps:
@@ III-D Augmented Robustness Ensemble
    $X_{T,i}\leftarrow\bar{R}^{-1/2}X_{T,i}.\quad i=1,...,N_{T}$ (2)
    After EA, the EEG trials are whitened, i.e., the average spatial covariance matrix becomes the identity matrix. Since EA is also performed in each source domain, EEG data distributions from different users become more consistent.
@@
    $\beta=0.05$ was used in our experiments. After data augmentation, we obtain the augmented calibration set $\mathcal{D}^{\prime}_{T}=\left\{\left(X_{T,i},y_{T,i}\right),\left(X^{\prime}_{T,i},y_{T,i}\right)\right\}_{i=1}^{N_{T}}$.
@@
    $\min_{\bm{\theta}}\mathbb{E}_{(X_{T},y_{T})\sim\mathcal{D}^{\prime}_{T}}\left[\max_{X^{adv}\in\mathcal{B}(X_{T},\epsilon)}\mathcal{L}\left(C_{\bm{\theta}}\left(X^{adv}_{T}\right),y_{T}\right)\right],$ (4)
@@
    $X^{adv,i}$ $=\text{Proj}_{X,\epsilon}\left(X^{adv,i-1}\right.$
    $\left.+\alpha\cdot\text{sign}(\nabla_{X^{adv,i-1}}\mathcal{L}({C_{\bm{\theta}}}(X^{adv,i-1}),y))\right),$ (6)
@@
    $\min_{\bm{\theta}}\mathbb{E}_{(X_{T},y_{T})\sim\mathcal{D}^{\prime}_{T},(\tilde{X}_{S},y_{S})\sim\tilde{\mathcal{D}}_{S}}$
    $\left[\max_{X^{adv}\in\mathcal{B}(X_{T},\epsilon)}\mathcal{L}\left(C_{\bm{\theta}}\left(X^{adv}_{T}\right),y_{T}\right)+\mathcal{L}\left(C_{\bm{\theta}}\left(\tilde{X}_{S}\right),y_{S}\right)\right].$ (7)
@@
The integration of data augmentation and adversarial training enables the classifier to learn a more generalizable classification boundary, as illustrated in Fig. 6.
@@ IV Experiment Settings
This section introduces our experiment settings, including the datasets, performance evaluation metrics, and baseline algorithms.
@@ IV-B Performance Evaluation Metrics
Fig. 7 shows a benign EEG sample, its PGD counterpart with perturbation magnitude $\epsilon=0.01$, and its noisy counterpart with perturbation magnitude $\eta=1$. The adversarial perturbation and the random noise differ significantly. Although the adversarial sample closely resembles the benign sample, it can still lead to a substantial drop in EEG classification performance.
@@ IV-C Baseline Algorithms
We compared our proposed ARE with the following 13 algorithms in the literature:
@@
For fair comparison, we also report the performance of ABAT ensemble (ABAT-E) and non-ensemble ARE (AR) in Section V. The robust training perturbation amplitude $\epsilon$ in ABAT, ABAT-E, AR and ARE were all set to 0.03 [27]. (1) and (2) were used in all experiments to align the source data and the target calibration data for improved performance. In the test phase, we aligned the test data with the Euclidean arithmetic mean $\bar{R}$ calculated on the target domain calibration set, to avoid data leakage.
@@ V-A Centralized Source-Free Transfer Learning Scenario
All source models were pre-trained in the centralized source-free transfer learning scenario, and then fine-tuned using the first 20%, 30%, 40%, 50% and 60% of the target user’s data as the calibration set. The average model accuracies are shown in Table II, and the detailed accuracies are shown in Fig. 8.
@@
Table II shows that the average accuracies of the four robust training approaches (ABAT, ABAT-E, AR, and ARE) on three types of samples (benign, adversarial, and noisy) were higher than the other 10 approaches across the three datasets. Additionally, their rank was consistently ABAT $<$ ABAT-E $<$ AR $<$ ARE, demonstrating the effectiveness of our proposed approaches.
Fig. 8 shows that:
@@
    As the amount of training data increased, the accuracies of the four robust training approaches (ABAT, ABAT-E, AR, and ARE) exhibited a clear upward trend across all sample types. In most cases, ARE achieved the highest classification accuracy, and AR the second highest, demonstrating again the effectiveness of our proposed approaches.
@@ V-B Baseline Performance in Centralized Source-Free Machine Learning
The classification accuracies of various baseline approaches on the benign samples are shown in the ‘Benign’ columns of Table II and the ‘Benign’ subfigures of Fig. 8. Domain adaptation approaches significantly outperformed models trained on the source or target data only. On average, domain adaptation approaches achieved a minimum accuracy of 64.85%, whereas the highest accuracy for models without adaptation was 56.63%. Most domain adaptation losses enhanced the classification accuracy on the benign target samples (with average accuracies over 66.26%), compared to the CE loss only, demonstrating the effectiveness of domain adaptation.
The classification accuracies of various baseline approaches on adversarial and noisy samples are shown in the ‘Adversarial’ and ‘Noisy’ columns of Table II, and the ‘Adversarial’ and ‘Noisy’ subfigures of Fig. 8. Observe that:
@@ V-C Federated Source-Free Transfer Learning Scenario
The first 20%, 30%, 40%, 50% and 60% of the target data were used as the calibration set in fine-tuning. All source models, except for the traditional CSP model, were pre-trained using federated source-free transfer learning. The average classification accuracies are shown in Table III, and the detailed accuracies are shown in Fig. 9.
Table III shows that, similar to in the centralized source-free transfer learning scenario, the average accuracies of the four robust training approaches (ABAT, ABAT-E, AR, and ARE) were always higher than the other 10 approaches, and the rank was consistently ABAT $<$ ABAT-E $<$ AR $<$ ARE. Fig. 9 shows that ARE was always the best-performing approach.
@@ V-D Source Data Perturbation Scenario
We combined the perturbed source data with the first 20%, 30%, 40%, 50% and 60% of the target data in training. The average classification accuracies are shown in Table IV, and the detailed accuracies in Fig. 10.
@@
    The four robust training approaches (ABAT, ABAT-E, AR, and ARE) again demonstrated better performance than others, and their rank was ABAT $\approx$ AR $<$ ABAT-E $\approx$ ARE. The utilization of source EEG data reduced performance difference between ABAT and AR; however, ensemble learning still achieved further performance improvement.
@@
The classification accuracies on the benign samples are shown in the ‘Benign’ columns of Table IV and the ‘Benign’ subfigures of Fig. 10. Compared to models trained only on source or target domain data, or on a mix of the two (S&T), domain adaptation algorithms notably improved the accuracies on benign target domain samples. Different domain adaptation losses did not significantly enhance accuracies over S+T on benign samples. This may be because mixing source and target data within each batch (S+T) enabled models to learn sufficient cross-domain knowledge, so further adaptation to the target data yielded limited benefit.
@@ V-E Comparison with Approaches without Privacy Protection
To further verify the performance of our proposed ARE, we combined the source domain data and the first 20%, 30%, 40%, 50% and 60% target domain data for model training, without imposing any privacy protection strategy. The average classification accuracies are shown in Table V, and the detailed accuracies are shown in Fig. 11. The four robust training approaches again achieved the highest accuracies.
@@ VI Conclusions
This paper has proposed an ARE approach, which leverages data alignment, data augmentation, adversarial training and ensemble learning to achieve more accurate and more robust EEG decoding. ARE is then integrated into three privacy protection scenarios (centralized source-free transfer, federated source-free transfer, and source data perturbation) to achieve simultaneously accurate classification, adversarial robustness, and privacy protection. Experiments on three public MI datasets demonstrated the effectiveness of ARE, even outperforming state-of-the-art transfer learning approaches that do not consider privacy protection at all. This is the first time that three major challenges in EEG-based BCIs (data scarcity and individual differences, adversarial vulnerability, and user privacy) can be accommodated simultaneously, significantly improving the practicalness of EEG decoding algorithms in real-world BCIs.

### v2

# A3E: Aligned and Augmented Adversarial Ensemble for Accurate, Robust and Privacy-Preserving EEG Decoding

@@ Abstract
An electroencephalogram (EEG) based brain-computer interface (BCI) enables direct communication between the brain and external devices. However, EEG-based BCIs face at least three major challenges in real-world applications: data scarcity and individual differences, adversarial vulnerability, and data privacy. While previous studies have addressed one or two of these issues, simultaneous accommodation of all three challenges remains challenging and unexplored. This paper fills this gap, by proposing an Aligned and Augmented Adversarial Ensemble (A3E) algorithm and integrating it into three privacy protection scenarios (centralized source-free transfer, federated source-free transfer, and source data perturbation), achieving simultaneously accurate decoding, adversarial robustness, and privacy protection of EEG-based BCIs. Experiments on three public EEG datasets demonstrated that our proposed approach outperformed over 10 classic and state-of-the-art approaches in both accuracy and robustness in all three privacy-preserving scenarios, even outperforming state-of-the-art transfer learning approaches that do not consider privacy protection at all. This is the first time that three major challenges in EEG-based BCIs can be addressed simultaneously, significantly improving the practicalness of EEG decoding in real-world BCIs.
@@ I Introduction
A typical closed-loop EEG-based BCI system includes four essential parts: signal acquisition, signal processing, machine learning, and a controller, as shown in Figure 1.
@@
However, to our knowledge, no study so far can address all three challenges simultaneously. This paper fills this gap. We demonstrate for the first time that it is feasible to achieve simultaneously accuracy, robust and privacy-preserving decoding of EEG-based BCIs through a single classifier. More specifically, we first propose aligned and augmented adversarial ensemble (A3E), which leverages data alignment, data augmentation, adversarial training, and ensemble learning to improve the classification accuracy and robustness, and then integrate it into three privacy-preserving transfer learning scenarios:
@@
    We propose A3E, which leverages data alignment, data augmentation, adversarial training, and ensemble learning to improve simultaneously the classification accuracy and adversarial robustness.
@@
    We integrate A3E with privacy-preserving transfer learning, achieving for the first time simultaneous accurate decoding, adversarial robustness, and privacy protection.
@@ II-C Privacy Protection
III Accurate, Robust, and Privacy-Preserving EEG Decoding
This section introduces our proposed approach for accurate, robust, and privacy-preserving EEG-Based BCIs, including the A3E algorithm and three privacy-preserving scenarios.
@@ III-B Overall Framework
In privacy-protection scenarios, the task model $\emph{C}_{\bm{\theta}}$ trained in the source domain or perturbed privacy-preserving source data $\tilde{\mathcal{D}}_{S}$ can be used to facilitate the training in the target domain using our proposed A3E, as shown in Figure 2 and Algorithm 1.
@@
Algorithm 1 Algorithm for accurate, robust and privacy-preserving EEG decoding.
@@ III-C1 Centralized Source-Free Transfer Learning
In this scenario, the source users trust each other, so they can share raw EEG data among themselves, but not with the target user. EEG data from multiple source users are aggregated to train a single source model, which is then utilized in transfer learning to train the target model, as illustrated in Figure 3. As the target user cannot access the source domain data, the source data privacy is protected.
@@ III-C2 Federated Source-Free Transfer Learning
In this more strict privacy-protection scenario, no source user wants to share his/her raw EEG data with other source users and the target user. A central server, without access to the source users’ private EEG data, maintains a global model and distributes it to each source user for updates. Each source user refines the global model parameters using his/her own EEG data and sends these updates back to the server for aggregation, as illustrated in Figure 4. Our recently proposed federated learning algorithm [44], which achieves superior performance compared to centralized training, is utilized in this paper.
@@ III-C3 Source Data Perturbation
A user-specific perturbation can be applied to each source domain to safeguard source user privacy while still leveraging the source data for target domain classifier training, as illustrated in Figure 5. This perturbation makes the user identity information unlearnable while preserving the BCI task related information. Our recently proposed privacy-preserving perturbation approach [26] is utilized in this paper.
@@
III-D Aligned and Augmented Adversarial Ensemble (A3E)
We propose A3E to train an accurate and robust EEG classifier for the target user, based on the target domain calibration set $\mathcal{D}_{T}=\left\{\left(X_{T,i},y_{T,i}\right)\right\}_{i=1}^{N_{T}}$. A3E has the following four steps:
@@ III-D Aligned and Augmented Adversarial Ensemble (A3E)
    $X_{T,i}\leftarrow\bar{R}^{-1/2}X_{T,i},\quad i=1,...,N_{T}.$ (2)
    After EA, the EEG trials are whitened, i.e., the average spatial covariance matrix becomes the identity matrix. Since EA is also performed in each source domain, EEG data distributions from different users become more consistent. Note that other alignment approaches, e.g., label alignment [46] and centroid alignment [47], may also be used.
@@
    $\beta=0.05$ was used in our experiments. After data augmentation, we obtain the augmented calibration set $\mathcal{D}^{\prime}_{T}=\left\{\left(X_{T,i},y_{T,i}\right),\left(X^{\prime}_{T,i},y_{T,i}\right)\right\}_{i=1}^{N_{T}}$. Note that other data augmentation approaches, e.g., channel reflection [49], may also be used.
@@
    $\min_{\bm{\theta}}\mathbb{E}_{(X_{T},y_{T})\sim\mathcal{D}^{\prime}_{T}}\left[\max_{X^{adv}_{T}\in\mathcal{B}(X_{T},\epsilon)}\mathcal{L}\left(C_{\bm{\theta}}\left(X^{adv}_{T}\right),y_{T}\right)\right],$ (4)
@@
    $X^{adv,i}=\text{Proj}_{X,\epsilon}\left(X^{adv,i-1}+\alpha\cdot\text{sign}(\nabla_{X^{adv,i-1}}\mathcal{L}({C_{\bm{\theta}}}(X^{adv,i-1}),y))\right),$ (6)
@@
    $\min_{\bm{\theta}}\mathbb{E}_{(X_{T},y_{T})\sim\mathcal{D}^{\prime}_{T},(\tilde{X}_{S},y_{S})\sim\tilde{\mathcal{D}}_{S}}\left[\max_{X^{adv}_{T}\in\mathcal{B}(X_{T},\epsilon)}\mathcal{L}\left(C_{\bm{\theta}}\left(X^{adv}_{T}\right),y_{T}\right)+\mathcal{L}\left(C_{\bm{\theta}}\left(\tilde{X}_{S}\right),y_{S}\right)\right].$ (7)
@@
The integration of data augmentation and adversarial training enables the classifier to learn a more generalizable classification boundary, as illustrated in Figure 6.
@@ IV Experiment Settings
This section introduces our experiment settings, including the datasets, performance evaluation metrics, and baseline algorithms. Our source code will be made available after this paper is accepted.
@@ IV-B Performance Evaluation Metrics
Figure 7 shows a benign EEG sample, its PGD counterpart with perturbation magnitude $\epsilon=0.01$, and its noisy counterpart with perturbation magnitude $\eta=1$. The adversarial perturbation and the random noise differ significantly. Although the adversarial sample closely resembles the benign sample, it can still lead to a substantial drop in EEG classification performance.
@@ IV-C Baseline Algorithms
We compared our proposed A3E with the following 13 algorithms in the literature:
@@
For fair comparison, we also report the performance of ABAT ensemble (ABAT-E) and non-ensemble A3E (A3) in Section V. The robust training perturbation amplitude $\epsilon$ in ABAT, ABAT-E, A3 and A3E were all set to 0.03 [27]. (1) and (2) were used in all experiments to align the source data and the target calibration data for improved performance. In the test phase, we aligned the test data with the Euclidean arithmetic mean $\bar{R}$ calculated on the target domain calibration set, to avoid data leakage.
@@ V-A Centralized Source-Free Transfer Learning Scenario
All source models were pre-trained in the centralized source-free transfer learning scenario, and then fine-tuned using the first 20%, 30%, 40%, 50% and 60% of the target user’s data as the calibration set. The average model accuracies are shown in Table II, and the detailed accuracies are shown in Figure 8.
@@
Table II shows that the average accuracies of the four robust training approaches (ABAT, ABAT-E, A3, and A3E) on three types of samples (benign, adversarial, and noisy) were higher than the other 10 approaches across the three datasets. Additionally, their rank was consistently ABAT $<$ ABAT-E $<$ A3 $<$ A3E, demonstrating the effectiveness of our proposed approaches.
Figure 8 shows that:
@@
    As the amount of training data increased, the accuracies of the four robust training approaches (ABAT, ABAT-E, A3, and A3E) exhibited a clear upward trend across all sample types. In most cases, A3E achieved the highest classification accuracy, and A3 the second highest, demonstrating again the effectiveness of our proposed approaches.
@@ V-B Baseline Performance in Centralized Source-Free Machine Learning
The classification accuracies of various baseline approaches on the benign samples are shown in the ‘Benign’ columns of Table II and the ‘Benign’ subfigures of Figure 8. Domain adaptation approaches significantly outperformed models trained on the source or target data only. On average, domain adaptation approaches achieved a minimum accuracy of 64.85%, whereas the highest accuracy for models without adaptation was 56.63%. Most domain adaptation losses enhanced the classification accuracy on the benign target samples (with average accuracies over 66.26%), compared to the CE loss only, demonstrating the effectiveness of domain adaptation.
The classification accuracies of various baseline approaches on adversarial and noisy samples are shown in the ‘Adversarial’ and ‘Noisy’ columns of Table II, and the ‘Adversarial’ and ‘Noisy’ subfigures of Figure 8. Observe that:
@@ V-C Federated Source-Free Transfer Learning Scenario
The first 20%, 30%, 40%, 50% and 60% of the target data were used as the calibration set in fine-tuning. All source models, except for the traditional CSP model, were pre-trained using federated source-free transfer learning. The average classification accuracies are shown in Table III, and the detailed accuracies are shown in Figure 9.
Table III shows that, similar to in the centralized source-free transfer learning scenario, the average accuracies of the four robust training approaches (ABAT, ABAT-E, A3, and A3E) were always higher than the other 10 approaches, and the rank was consistently ABAT $<$ ABAT-E $<$ A3 $<$ A3E. Figure 9 shows that A3E was always the best-performing approach.
@@ V-D Source Data Perturbation Scenario
We combined the perturbed source data with the first 20%, 30%, 40%, 50% and 60% of the target data in training. The average classification accuracies are shown in Table IV, and the detailed accuracies in Figure 10.
@@
    The four robust training approaches (ABAT, ABAT-E, A3, and A3E) again demonstrated better performance than others, and their rank was ABAT $\approx$ A3 $<$ ABAT-E $\approx$ A3E. The utilization of source EEG data reduced performance difference between ABAT and A3; however, ensemble learning still achieved further performance improvement.
@@
The classification accuracies on the benign samples are shown in the ‘Benign’ columns of Table IV and the ‘Benign’ subfigures of Figure 10. Compared to models trained only on source or target domain data, or on a mix of the two (S&T), domain adaptation algorithms notably improved the accuracies on benign target domain samples. Different domain adaptation losses did not significantly enhance accuracies over S+T on benign samples. This may be because mixing source and target data within each batch (S+T) enabled models to learn sufficient cross-domain knowledge, so further adaptation to the target data yielded limited benefit.
@@ V-E Comparison with Approaches without Privacy Protection
To further verify the performance of our proposed A3E, we combined the source domain data and the first 20%, 30%, 40%, 50% and 60% target domain data for model training, without imposing any privacy protection strategy. The average classification accuracies are shown in Table V, and the detailed accuracies are shown in Figure 11. The four robust training approaches again achieved the highest accuracies.
@@ VI Conclusions
This paper has proposed an A3E approach, which leverages data alignment, data augmentation, adversarial training and ensemble learning to achieve more accurate and more robust EEG decoding. A3E is then integrated into three privacy protection scenarios (centralized source-free transfer, federated source-free transfer, and source data perturbation) to achieve simultaneously accurate classification, adversarial robustness, and privacy protection. Experiments on three public MI datasets demonstrated the effectiveness of A3E, even outperforming state-of-the-art transfer learning approaches that do not consider privacy protection at all. This is the first time that three major challenges in EEG-based BCIs (data scarcity and individual differences, adversarial vulnerability, and user privacy) can be accommodated simultaneously, significantly improving the practicalness of EEG decoding algorithms in real-world BCIs.